OpenStudy (anonymous):

Fan & Medal A floor plan is given for the first floor of a new house. One inch represents 6 feet. If a small closet is 1/4 in by 1/3 in on the floor plan, what are the actual dimensions? 1.36 ft by 1.5 ft 1.75 ft by 2 ft 1.5 ft by 2 ft 2 ft by 2 ft

OpenStudy (anonymous):

c

OpenStudy (anonymous):

An inch on the floor plan is 6 feet in real life. So 1/4 inch = 6/4 feet = 1.5 feet 1/3 inch = 6/3 feet = 2 feet

OpenStudy (anonymous):

You get it ?

OpenStudy (anonymous):

yes thank you

OpenStudy (danjs):

1 in = 6 ft \[\frac{ 1 ~inch}{ 6~ft } = 1\] so you can multiply your dimensions by that ratio, same as multiplying by 1, so the units work out correctly

OpenStudy (danjs):

\[\frac{ 1/4~inch }{ 1 } *\frac{ 6~ft }{ 1~inch} = \]

OpenStudy (danjs):

inches cancel, left with ft